# Q-learning with Posterior Sampling

### v2

@@ 1 Introduction
In a Reinforcement Learning (RL) problem, an agent interacts sequentially with an unknown environment and uses the observed outcomes to learn an interaction strategy. The underlying mathematical model for RL is a Markov Decision Process (MDP). In the tabular episodic RL setting, there is an underlying MDP with a finite state space $\cal S$}, a finite action space $\cal A$} and a planning horizon $H$. On taking an action $a$ in state $s$ at step $h$, the environment produces a reward and next state from the (unknown) reward model $R_{h}(s,a)$ and transition probability model $P_{h}(s,a)$ of the underlying MDP.
Q-learning Watkins and Dayan (1992) is a classic dynamic programming (DP)-based algorithm for RL. The DP equations (aka Bellman equations) provide a recursive expression for the optimal expected reward achievable from any state and action of the MDP, aka the $Q$-values, in terms of the optimal value achievable in the next state. Specifically, for any given $s\in{\cal S},a\in{\cal A},h\in[H]$, the Q-value $Q_{h}(s,a)$ is given by:
@@
$V_{h+1}(s^{\prime})\mathrel{\mathop{\ordinarycolon}}=\max_{a^{\prime}\in{\cal A}}Q_{h+1}(s^{\prime},a^{\prime}).$
Here, the terminal value $V_{H+1}(s)$ is 0 for all $s$. The optimal action in state $s$ is then given by the argmax action in the above.
In the reinforcement learning setting, where the reward and transition models of the MDP are unknown, the $Q$-learning algorithm uses the celebrated Temporal Difference (TD) learning idea Sutton (1988) to construct increasingly accurate estimates of $Q$-values using past observations. The key idea here is to construct an estimate of the right hand side of the Bellman equation, aka target, by bootstrapping the current estimate $\widehat{V}_{h+1}$ for the next step value function. That is, on playing an action $a$ in state $s$ at step $h$, and observing reward $r_{h}$ and next state $s^{\prime}$, the target $z$ is typically constructed as:
$z\mathrel{\mathop{\ordinarycolon}}=r_{h}+\widehat{V}_{h+1}(s^{\prime})$
And then the estimate ${\widehat{Q}}_{h}(s,a)$ for $Q_{h}(s,a)$ is updated to fit the Bellman equations using the Q-learning update rule [^2] Or, $\widehat{Q}_{h}(s,a)\leftarrow\widehat{Q}_{h}(s,a)+\alpha_{n}(z-\widehat{Q}_{h}(s,a))$ where $z-\widehat{Q}_{h}(s,a)$ is called the Temporal Difference (TD).
@@
There are several ways to interpret the Q-learning update rule. The traditional frequentist interpretation popularized by Mnih et al. (2015) interprets this update as a gradient descent step for a least squares regression error minimization problem. Another, arguably more insightful, interpretation of $Q$-learning can be obtained using Bayesian inference theory. Specifically, if we assume a Gaussian prior
@@
(See Section 3.1 for further details.)
Importantly, the Bayesian posterior tracks not just the mean but also the variance or uncertainty in the $Q$-value estimate. This uncertainty quantification is particularly useful for managing the exploration-exploitation tradeoff for regret minimization in reinforcement learning.
Figure 1: A comparison of PSQL(a heuristic derived from PSQL), UCBQL Jin et al. (2018); (Staged) RandQLTiapkin et al. (2023), and RLSVI Russo (2019) in a chain MDP environment (for details, and more experiments, see Appendix A).
Intuitively, the state and actions with a small number of past visits (i.e., small $n$) have large uncertainty in their current $Q$-value estimate, and should be explored more. The posterior sampling approaches implement this idea by simply taking a sample from the posterior, which is likely to be closer to the mean (less exploration) for actions with small posterior variance, and away from the mean (more exploration) for those with large variance.
@@
Popularized by their success in the multi-armed bandit settings Thompson (1933); Chapelle and Li (2011); Kaufmann et al. (2012); Agrawal and Goyal (2017), and in deep reinforcement learning regimes (Osband et al., 2016a; Fortunato et al., 2017; Azizzadenesheli et al., 2018; Li et al., 2021b; Fan and Ming, 2021; Sasso et al., 2023), the posterior sampling approaches are generally believed to be more efficient in managing the exploration-exploitation tradeoff than their UCB (Upper Confidence Bound) counterparts. Our preliminary experiments (see Figure 1) suggest that this is also the case for our Q-learning approach in the tabular RL setting.  [^3] For the empirical study reported here, we implement a vanilla version of posterior sampling with Q-learning. The PSQL algorithm presented later modifies the target computation slightly as described later in Section 3, for the sake of theoretical analysis. However, obtaining provable guarantees for posterior sampling approaches have historically been more challenging.
@@
Another closely related approach with provable regret bounds is the RLSVI (Randomized Least Squared Value Iteration) algorithm by Osband et al. (2016b, 2019); Russo (2019); Zanette et al. (2020); Agrawal et al. (2021); Xiong et al. (2022). The RLSVI algorithm is an approximate value iteration-based approach that can be interpreted as maintaining “empirical posteriors" over the value functions by injecting noise. However, in the tabular setting considered here, RLSVI reduces to a model-based algorithm. (See Section 2.1 for further comparison.) There are several other model-based posterior sampling algorithms in the literature with near optimal regret bounds Osband et al. (2013); Osband and Van Roy (2017); Ouyang et al. (2017); Agrawal and Jia (2022); Agarwal and Zhang (2022); Tiapkin et al. (2022). Model-based algorithms directly estimate the reward and/or transition model, instead of the implied optimal value functions, or policy parameters. In many settings, model-based algorithms can be more sample efficient. But, model-free approaches like Q-learning have gained popularity in practice because of their simplicity and flexibility, and underlie most successful modern deep RL algorithms (e.g., DQN (Mnih et al., 2013), DDQN (van Hasselt et al., 2015), A3C (Mnih et al., 2016)). Provable regret bounds for a simple posterior sampling based Q-learning algorithm like PSQL, therefore, still remains a problem of significant interest.
@@
  • We propose the Q-learning with Posterior Sampling (PSQL) algorithm that combines the simplicity and flexibility of the popular Q-learning algorithm with efficient exploration provided by the Bayesian posterior sampling approach. Our preliminary experiments demonstrate promising empirical performance of this simple algorithm compared to contemporary approaches. (See Section 3 for algorithm design and Appendix A for experiments.)
  • We provide a novel derivation of Q-learning as a solution to a Bayesian inference problem with a regularized Evidence Lower Bound (ELBO) objective. Besides forming the basis of our PSQL algorithm design, this derivation provides a more insightful interpretation of the modified learning rates introduced in some previous works on $Q$-learning (e.g., (Jin et al., 2018)) to obtain provable regret bounds. (See Section 3.1.)
  • We prove a near-optimal regret bound of $\tilde{O}(H^{2}\sqrt{SAT})$ for PSQL which closely match the known lower bound of $\Omega(H\sqrt{SAT})$ (Jin et al., 2018). Our result improves the regret bounds available for the closely related approach of RLSVI Russo (2019) and matches those recently derived by Tiapkin et al. (2023) for a more complex posterior sampling based algorithm Staged-RandQL. (See Section 2,4.)
@@ 2 Our setting and main result
We consider an online reinforcement learning setting. In this setting, the algorithm interacts with environment in $K$ sequential episodes, each containing $H$ steps. At step $h=1,\ldots,H$ of each episode $k$, the algorithm observes the current state $s_{k,h}$, takes an action $a_{k,h}$ and observes a reward $r_{k,h}$ and the next state $s_{k,h+1}$.
The reward and next state are generated by the environment according to a fixed underlying MDP $({\cal S},{\cal A},R,P)$, so that $\Pr(s_{k,h+1}=s^{\prime}|s_{k,h}=s,a_{k,h}=a)=P_{h}(s,a,s^{\prime}),\mathbb{E}[r_{k,h}|s_{k,h}=s,a_{k,h}=a]=R_{h}(s,a)$. However, the reward functions and the transition probability distributions $R_{h},P_{h},h=1,\ldots,H$ are apriori unknown to the algorithm. The goal is to minimize total regret compared to the optimal value given by the dynamic programming equation (1). Specifically, let $\pi_{k}$ denote the policy used by the algorithm in episode $k$, so that $a_{k,h}=\pi^{k}(s_{k,h})$. We aim to bound regret, defined as
@@
A regret minimizing algorithm uses past observations to learn about the MDP and improve its actions. Since the algorithm can only observe the environment’s response at a visited state and action, a main challenge in this problem is managing the exploration-exploitation tradeoff. This refers to the dilemma between picking actions that are most likely to be optimal according to the observations made so far, versus picking actions for that allow visiting less-explored states and actions. Following work on managing exploration-exploitation tradeoffs in multi-armed bandit (MAB) settings, two main approaches for exploration exploitation are optimistic approaches based on UCB and posterior sampling approaches (aka Thompson Sampling in MAB).
@@
Next, we derive the posterior sampling approach considered in this paper, and relate it to the other popular versions of Q-learning for regret minimization in RL.
@@ 2.1 Related works
UCB UCBQL Jin et al. (2018) $\tilde{O}(H^{1.5}\sqrt{SAT})$ Q-learning with UCB
Q-EarlySettled-Advantage Li et al. (2021a) $\tilde{O}(H\sqrt{SAT})$ Q-learning with UCB
@@
Posterior sampling Conditional Posterior Sampling Dann et al. (2021) $\tilde{O}(HSA\sqrt{T})$ computationally intractable
RLSVI Russo (2019) $\tilde{O}(H^{3}S^{1.5}\sqrt{AT})$ approximate value iteration with noise
C-RLSVI Agrawal et al. (2021) $\tilde{O}(H^{2}S\sqrt{AT})$ approximate value iteration with noise
@@
Staged RandQL Tiapkin et al. (2023) $\tilde{O}(H^{2}\sqrt{SAT})$ Q-learning with learning-rate randomization
@@
PSQL [this work] $\tilde{O}(H^{2}\sqrt{SAT})$ Q-learning with Gaussian posteriors
Lower bound Jin et al. (2018) $\Omega(H\sqrt{SAT})$ -
@@
Table 1: Comparison of our regret bound to related wors ( Dann et al. (2021) is in function approximation setting).
@@
Our work falls under the umbrella of online episodic reinforcement learning, specifically on regret minimization in tabular setting. In the category of the Upper Confidence Bound (UCB)-based algorithms, there is a huge body of research both on model-based (Bartlett and Tewari, 2012; Azar et al., 2017; Fruit et al., 2018; Zanette and Brunskill, 2019; Zhang et al., 2020; Boone and Zhang, 2024), and model-free Jin et al. (2018); Bai et al. (2019); Ménard et al. (2021); Zhang and Xie (2023); Agrawal and Agrawal (2024) algorithms. In-fact, Jin et al. (2018) were the first to provide a near-optimal worst-case regret bound of $\tilde{O}(\sqrt{H^{3}SAT})$, subsequently improved to $\tilde{O}(H\sqrt{SAT})$ by Zhang et al. (2020); Li et al. (2021a).
@@
Motivated by the superior empirical performance of Bayesian posterior sampling approaches compared to their UCB counterparts (Chapelle and Li, 2011; Kaufmann et al., 2012; Osband et al., 2013; Osband and Van Roy, 2017; Osband et al., 2019) there have been several attempts at deriving provable regret bounds for these approaches in the episodic RL setting. Among model-based approaches, near optimal regret bounds have been established for approaches that use (typically Dirichlet) posteriors on transition models Ouyang et al. (2017); Agrawal and Jia (2017); Tiapkin et al. (2022). There have been relatively limited studies on model-free, sample-efficient and computationally efficient Bayesian algorithms. Dann et al. (2021) proposed one such framework but is computationally intractable. Our work aims to fill this gap.
@@
A popular approach closely related to posterior sampling is Randomized Least Square Value Iteration (RLSVI) (Osband et al., 2016b, 2019; Russo, 2019; Zanette et al., 2020). In RLSVI, the exploration is carried out by injecting randomized uncorrelated noise to the reward samples, followed by a re-fitting of a $Q$-function estimate by solving a least squares problem on all the past data, incurring heavy computation and storage costs. This process has been interpreted as forming an approximate posterior distribution over value functions. RLSVI too enjoys a “superior-than-UCB” empirical performance. However, its worst-case regret bounds Russo (2019) remain suboptimal (see Table 1) in their dependence on the size of the state space.
@@
More recently, Tiapkin et al. (2023) proposed (RandQL and Staged-RandQL) algorithms that are model-free, tractable and enjoy $\tilde{O}(H^{2}\sqrt{SAT})$ regret by randomizing the learning rates of the Q-learning update rule. Their algorithmic design is based on Dirichlet posteriors on transition models and efficient implmentaton of the implied distribution on $Q$-value estimates via learning rate randomization. In our preliminary experiments (see Figure 1 and Section A), our PSQL approach with simple Gaussian based posteriors shows significantly better performance compared to these algorithms.

### v3

@@ 1 Introduction
In an online Reinforcement Learning (RL) problem, an agent interacts sequentially with an unknown environment and uses the observed outcomes to learn an interaction strategy. The underlying mathematical model for RL is a Markov Decision Process (MDP). In the tabular episodic setting, the MDP has a finite state space $\cal S$}, a finite action space $\cal A$} and a planning horizon $H$. On taking an action $a$ in state $s$ at step $h$, the environment produces a reward and next state from the (unknown) reward model $R_{h}(s,a)$ and transition probability model $P_{h}(s,a)$ of the underlying MDP.
Q-learning (Watkins and Dayan, 1992) is a classic dynamic programming (DP)-based algorithm for RL. The DP equations (aka Bellman equations) provide a recursive expression for the optimal expected reward achievable from any state and action of the MDP, aka the $Q$-values, in terms of the optimal value achievable in the next state. Specifically, for any given $s\in{\cal S},a\in{\cal A},h\in[H]$, the Q-value $Q_{h}(s,a)$ is given by:
@@
$V_{h+1}(s^{\prime})\mathrel{\mathop{\ordinarycolon}}=\max_{a^{\prime}\in{\cal A}}Q_{h+1}(s^{\prime},a^{\prime}),$
with $V_{H+1}(s)=0,\,\forall\,s$. The optimal action in state $s$ is then given by the argmax action in the above.
When the reward and transition models of the MDP are unknown, the $Q$-learning algorithm uses the celebrated Temporal Difference (TD) learning idea (Sutton, 1988) to construct increasingly accurate estimates of $Q$-values using past observations. The key idea here is to construct an estimate of the right hand side of the Bellman equation, aka target, by bootstrapping the current estimate $\widehat{V}_{h+1}$ for the next step value function. That is, on playing an action $a$ in state $s$ at step $h$, and observing reward $r_{h}$ and next state $s^{\prime}$, the target $z$ is typically constructed as: $z\mathrel{\mathop{\ordinarycolon}}=r_{h}+\widehat{V}_{h+1}(s^{\prime})$.
And the estimate ${\widehat{Q}}_{h}(s,a)$ for $Q_{h}(s,a)$ is updated to fit the Bellman equations using the Q-learning update rule [^2] Or, $\widehat{Q}_{h}(s,a)\leftarrow\widehat{Q}_{h}(s,a)+\alpha_{n}(z-\widehat{Q}_{h}(s,a))$ where $z-\widehat{Q}_{h}(s,a)$ is called the Temporal Difference (TD).
@@
There are several ways to interpret the Q-learning update rule. The traditional frequentist interpretation popularized by Mnih et al. (2015) interprets this update as a gradient descent step for a least squares regression error minimization problem. We propose a more insightful interpretation of $Q$-learning obtained using Bayesian inference theory (details in Section 3.1). Specifically, if we assume a Gaussian prior
@@
Importantly, the Bayesian posterior tracks not just the mean but also the variance or uncertainty in the $Q$-value estimate. Intuitively, the state and actions with a small number of past visits (i.e., small $n$) have large uncertainty in their current $Q$-value estimate, and should be explored more. The posterior sampling approaches implement this idea by simply taking a sample from the posterior, which is likely to be closer to the mean (less exploration) for actions with small posterior variance, and away from the mean (more exploration) for those with large variance. This uncertainty quantification is useful for managing the exploration-exploitation tradeoff for regret minimization. The exploration methodology is distinct from algorithms that use additive bonuses or randomized perturbations in the estimates.
Figure 1: Performance comparison of PSQL(a heuristic derived from PSQL), UCBQL (Jin et al., 2018); Staged-RandQL(Tiapkin et al., 2023), and RLSVI (Russo, 2019) in a chain MDP environment (for details, and more experiments, see Appendix A).
@@
Popularized by their success in the multi-armed bandit settings (Thompson, 1933; Chapelle and Li, 2011; Kaufmann et al., 2012; Agrawal and Goyal, 2017), and in deep reinforcement learning regimes (Osband et al., 2016a; Fortunato et al., 2017; Azizzadenesheli et al., 2018; Li et al., 2021b; Fan and Ming, 2021; Sasso et al., 2023), the posterior sampling approaches are generally believed to be more efficient in managing the exploration-exploitation tradeoff than their UCB (Upper Confidence Bound) counterparts. Our preliminary experiments (see Figure 1) suggest that this is also the case for our Q-learning approach in the tabular RL setting.  [^3] For the empirical study reported here, we implement a vanilla version of posterior sampling with Q-learning. The PSQL algorithm presented later modifies the target computation as described later in Section 3, for the sake of theoretical analysis. However, obtaining provable guarantees for posterior sampling approaches have historically been more challenging.
@@
Another closely related approach with provable regret bounds is the RLSVI (Randomized Least Squared Value Iteration) algorithm by Osband et al. (2016b, 2019); Russo (2019); Zanette et al. (2020); Agrawal et al. (2021); Xiong et al. (2022). The RLSVI algorithm is an approximate value iteration-based approach that can be interpreted as maintaining “empirical posteriors" over the value functions by injecting noise. However, in the tabular setting considered here, RLSVI reduces to a model-based algorithm. (See Section 2.1 for further comparisons.) There are several other model-based posterior sampling algorithms in the literature with near optimal regret bounds (Osband et al., 2013; Osband and Van Roy, 2017; Ouyang et al., 2017; Agrawal and Jia, 2022; Agarwal and Zhang, 2022; Tiapkin et al., 2022). Model-based algorithms directly estimate the reward and/or transition model, instead of the implied optimal value functions, or policy parameters. In many settings, model-based algorithms can be more sample efficient. But, model-free approaches like Q-learning have gained popularity in practice because of their simplicity and flexibility, and underlie most successful modern deep RL algorithms (e.g., DQN Mnih et al. (2013), DDQN van Hasselt et al. (2015), A3C Mnih et al. (2016)). Provable regret bounds for a simple posterior sampling based Q-learning algorithm like PSQL, therefore, still remains a problem of significant interest.
@@
  • We propose the Q-learning with Posterior Sampling (PSQL) algorithm that is the first Q-learning algorithm with natural and efficient exploration provided by the Bayesian posterior sampling approach. Our preliminary experiments demonstrate promising empirical performance of this simple algorithm compared to contemporary approaches. (See Section 3 for algorithm design and Appendix A for experiments.)
  • We provide a novel derivation of Q-learning as a solution to a Bayesian inference problem with a regularized Evidence Lower Bound (ELBO) objective. Besides forming the basis of our PSQL algorithm design, this derivation provides a more insightful interpretation of the learning rates introduced in some previous works on $Q$-learning (e.g., Jin et al. (2018)) to obtain provable regret bounds. (See Section 3.1.)
  • We prove a near-optimal regret bound of $\tilde{O}(H^{2}\sqrt{SAT})$ for PSQL which closely match the known lower bound of $\Omega(H\sqrt{SAT})$ (Jin et al., 2018). Our result improves the regret bounds available for the closely related approach of RLSVI (Russo, 2019) and matches those recently derived by Tiapkin et al. (2023) for a more complex posterior sampling based algorithm Staged-RandQL. (See Section 24.)
@@ 2 Our setting and main result
In the online reinforcement learning setting, the algorithm interacts with environment in $K$ sequential episodes, each containing $H$ steps. At step $h=1,\ldots,H$ of each episode $k$, the algorithm observes the current state $s_{k,h}$, takes an action $a_{k,h}$ and observes a reward $r_{k,h}$ and the next state $s_{k,h+1}$.
The reward and next state are generated by the environment according to a fixed underlying MDP $({\cal S},{\cal A},R,P)$, so that $\Pr(s_{k,h+1}=s^{\prime}|s_{k,h}=s,a_{k,h}=a)=P_{h}(s,a,s^{\prime}),\mathbb{E}[r_{k,h}|s_{k,h}=s,a_{k,h}=a]=R_{h}(s,a)$. However, the reward functions and the transition probability distributions $R_{h},P_{h},h=1,\ldots,H$ are apriori unknown to the algorithm. The goal is to minimize total regret compared to the optimal value given by the dynamic programming equation (1). Specifically, let $\pi_{k}$ denote the policy used by the algorithm in episode $k$, so that $a_{k,h}=\pi_{k}(s_{k,h})$. We aim to bound regret, defined as
@@
Since the algorithm can only observe the environment’s response at a visited state and action, a main challenge in this problem is managing the exploration-exploitation tradeoff. This refers to the dilemma between picking actions that are most likely to be optimal according to the observations made so far, versus picking actions for that allow visiting less-explored states and actions. The two main approaches for managing exploration-exploitation tradeoffs are the optimistic approaches based on UCB and posterior sampling approaches (aka Thompson Sampling in multi-armed bandit settings).
@@ 2.1 Related works
UCB UCBQL (Jin et al., 2018) $\tilde{O}(H^{1.5}\sqrt{SAT})$ Q-learning with UCB
Q-EarlySettled-Advantage (Li et al., 2021a) $\tilde{O}(H\sqrt{SAT})$ Q-learning with UCB
@@
Posterior sampl. Conditional Posterior Sampling (Dann et al., 2021) $\tilde{O}(HSA\sqrt{T})$ computationally intractable
RLSVI (Russo, 2019) $\tilde{O}(H^{3}S^{1.5}\sqrt{AT})$ approximate value iteration
C-RLSVI (Agrawal et al., 2021) $\tilde{O}(H^{2}S\sqrt{AT})$ approximate value iteration
@@
Staged RandQL (Tiapkin et al., 2023) $\tilde{O}(H^{2}\sqrt{SAT})$ randomized learning-rates
@@
PSQL [this work] $\tilde{O}(H^{2}\sqrt{SAT})$ Gaussian posteriors on Q-values
Lower bound (Jin et al., 2018) $\Omega(H\sqrt{SAT})$ -
@@
Table 1: Comparison of our regret bound to related works ( Dann et al. (2021) is in function approximation setting).
@@
Our work falls under the umbrella of online episodic reinforcement learning on regret minimization in tabular setting. In the category of the Upper Confidence Bound (UCB)-based algorithms, there is a huge body of research both on model-based (Bartlett and Tewari, 2012; Azar et al., 2017; Fruit et al., 2018; Zanette and Brunskill, 2019; Zhang et al., 2020; Boone and Zhang, 2024), and model-free (Jin et al., 2018; Bai et al., 2019; Ménard et al., 2021; Zhang and Xie, 2023; Agrawal and Agrawal, 2024) algorithms. In-fact, Jin et al. (2018) were the first to provide a near-optimal worst-case regret bound of $\tilde{O}(\sqrt{H^{3}SAT})$, subsequently improved to $\tilde{O}(H\sqrt{SAT})$ by Zhang et al. (2020); Li et al. (2021a).
@@
Motivated by the superior empirical performance of Bayesian posterior sampling approaches compared to their UCB counterparts (Chapelle and Li, 2011; Kaufmann et al., 2012; Osband et al., 2013; Osband and Van Roy, 2017; Osband et al., 2019) there have been several attempts at deriving provable regret bounds for these approaches in the episodic RL setting. Among model-based approaches, near optimal regret bounds have been established for approaches that use (typically Dirichlet) posteriors on transition models (Ouyang et al., 2017; Agrawal and Jia, 2017; Tiapkin et al., 2022). There have been relatively limited studies on model-free, sample-efficient and computationally efficient Bayesian algorithms. Dann et al. (2021) proposed one such framework but is computationally intractable. Our work aims to fill this gap.
@@
A popular approach closely related to posterior sampling is Randomized Least Square Value Iteration (RLSVI) (Osband et al., 2016b, 2019; Russo, 2019; Zanette et al., 2020). In RLSVI, the exploration is carried out by injecting randomized uncorrelated noise to the reward samples, followed by a re-fitting of a $Q$-function estimate by solving a least squares problem on all the past data, incurring heavy computation and storage costs. This process has been interpreted as forming an approximate posterior distribution over value functions. RLSVI too enjoys a “superior-than-UCB” empirical performance. However, its worst-case regret bounds (Russo, 2019) remain suboptimal (see Table 1) in their dependence on the size of the state space.
@@
More recently, Tiapkin et al. (2023) proposed (RandQL and Staged-RandQL) algorithms that are model-free, tractable and enjoy $\tilde{O}(H^{2}\sqrt{SAT})$ regret by randomizing the learning rates of the Q-learning update rule. Their algorithmic design is based on Dirichlet posteriors on transition models and efficient implementation of the implied distribution on $Q$-value estimates via learning rate randomization. Our algorithm is much simpler with far lesser randomized sampling steps and in our preliminary experiments (see Figure 1 and Section A), our PSQL approach with simple Gaussian based posteriors shows better/comparable performance compared to these algorithms.